OpenStudy (anonymous):

You throw a water balloon downward from the roof of a building at a speed of 3.5 m/s. How fast will the balloon be moving when it hits the sidewalk 8.0 m below? a. 13 m/s c. 160 m/s b. 12.7 m/s d. 169 m/s

OpenStudy (anonymous):

@Fifciol

OpenStudy (anonymous):

Please and thank you.

OpenStudy (fifciol):

Apply the conservation of mechanical energy: sum of kinetic and potential energy must remain the same. You are free to choose your potential energy level to be zero. I prefer to choose the sidewalk level as zero to simplify calculations. Can you write down what is the total energy of system at level 8m above sidewalk?

OpenStudy (anonymous):

Well what i put on there is all the information that is given to me. I feel like this online class is meaning to fail me XD @Fifciol

OpenStudy (fifciol):

You should really start to learn physics if you want to solve these very simple problems and pass the course. If you want the exact answers rather than tutoring this is not the site you are looking for. But I'll write the energy equation if this could help you: $$\frac{mv_0^2}{2}+mgh=\frac{mv^2}{2}$$solve for v

OpenStudy (anonymous):

I take an online class that does not give me any information beforehand @Fifciol It's ridiculous. A complete downer if you ask me.

OpenStudy (issy14):

Here yo go: Equation Vf ( Final Velocity) = Square root of Vo^2 ( initial velocity) + 2(9.81)(8m) = 12.97m/s round it up to 13m/s. Vf= Square root of Vo^2 + ad